OpenStudy (anonymous):

What is the volume of a cone with the given dimensions? Radius = 4 cm ; Height = 10 cm Round the answer to the nearest hundredth. Use 3.14 for pi.

OpenStudy (anonymous):

\[V=\pi r ^{2}\frac{ h }{ 3 }\]\[r ^{2}=16\]\[\frac{ h }{ 3 }=3.333...\]------------\[V=3.14\times16\times3.333=167.467\]

OpenStudy (anonymous):

volume\[=\frac{ 1 }{ 3 }\pi r^2h\]

OpenStudy (anonymous):

Do you need help with anything else @kurtisjb?

OpenStudy (anonymous):

...and offline -_-